OpenStudy (anonymous):

What is the simplified form of x minus 3 over x squared plus x minus 12 • x plus 4 over x squared plus 8x plus 16 ?

OpenStudy (saifoo.khan):

\[\frac{x-3}{x^2+x-12} \times \frac{x+4}{x^2+8x+16}\]

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

factorize.

OpenStudy (saifoo.khan):

\[\frac{x-3}{(x-3)(x+4)} \times \frac{(x+4)}{(x+4)(x+4)}\]

OpenStudy (saifoo.khan):

can you now solve now:?

OpenStudy (anonymous):

1/(x-3)(x+4)

OpenStudy (anonymous):

?